OpenStudy (amistre64):

in a RxRxRxR; what is the volume of rotation of the equation: y = 2x+7z^2-4w^2+9; from 0 to 2 in all directions. Rotated about the y axis?

OpenStudy (anonymous):

what is a RxRxRxR? a hypercube?

OpenStudy (amistre64):

its a hyperspace, yes. if thats the proper term :)

OpenStudy (anonymous):

y=14z^(2)-8w^(2)+2x+9, I am not sure on this answer but looks right.......

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

hyperspace? Wow, Star Trek is getting closer to reality. try using integration

OpenStudy (amistre64):

that looks closer to a derivative than a volume of revolution.

OpenStudy (anonymous):

I tried

OpenStudy (amistre64):

you did :) and thats good

OpenStudy (amistre64):

im curious if i would use the shell method or not :)

OpenStudy (anonymous):

I like the plate method

OpenStudy (amistre64):

so what is the equation for volume in hyperspace? im sure its the integration of volume; since area ups to volume

OpenStudy (anonymous):

it's.... a double integral?

OpenStudy (amistre64):

\[\int 2pi\ r\ dr=pi\ r^2\] \[\int pi\ r^2\ dr=\frac{pi}{3}\ r^3\]

OpenStudy (amistre64):

what it is? i got no idea yet :)